OpenStudy (anonymous):

how do you solve this system of linear equations? x+y=240 18x+12y+3120

OpenStudy (anonymous):

=3120*

OpenStudy (anonymous):

deja vu

OpenStudy (amistre64):

id assume the way youd solve it as you do any system of equations

OpenStudy (anonymous):

but i dont know how to do any systems of equations...

OpenStudy (anonymous):

you should get \[x = 40,\ \ y = 200\]

OpenStudy (amistre64):

then i would advise you to go back and review the section that teaches you those methods before trying to learn how to actually apply those methods

OpenStudy (anonymous):

You have 3 options to solving systems of linear equations (with 2 variables x and y): graphing substitution elimination

OpenStudy (anonymous):

but, i learned this in algebra 1, and now im in algebra 2 and forgot

OpenStudy (amistre64):

thats a contradiction; if you learned it, then you would know it :)

OpenStudy (amistre64):

there are many methods to solving systems of equations; some more labor intensive than others

OpenStudy (amistre64):

the main objective to understand is that a solution means that when x and y have solid values; that they will produce the desired results

OpenStudy (anonymous):

but it was in 7th grade now im in 9th and i need one of which i could show work.

OpenStudy (amistre64):

these equation can be graphed by 2 lines; and where they cross is where the solution to the system is

OpenStudy (anonymous):

so where they cross will be an ordered pair tho

OpenStudy (amistre64):

yes, x = ? and y=? :)

OpenStudy (anonymous):

x = 40 and y = 200

OpenStudy (amistre64):

i would go after elimination; which is simply multiplying the top by -18 and adding the 2 together

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhh.

OpenStudy (amistre64):

x + y =240 ; *-18 18x+12y =3120 -18x -18y =240(-18) 18x +12y =3120 ------------------- -6y = 3120 - 240(-18) y = (3120 - 240(-18))/-6

OpenStudy (amistre64):

i get y = -1240 that way tho

OpenStudy (amistre64):

oh, i subtracted the right side instead of adding lol

OpenStudy (amistre64):

y = (3120 + 240(-18))/-6 = 200 :)

OpenStudy (anonymous):

thanks